The population of the world in 1995 was 5.7 billion people. The observed population growth rate was 2% per year.
-Assuming the growth rate doesn't change, what will the population be in the year 2046?
-In what year will the worlds population have tripled?
-The predicted sustainable population of the earth is 20.1 billion people. When will Earth have too many people?
I got 15.6 billion for part a and I need help on parts b and c. Thank you!!

Answer by stanbon(75887)

The population of the world in 1995 was 5.7 billion people. The observed population growth rate was 2% per year.
-Assuming the growth rate doesn't change, what will the population be in the year 2046?
1995 to 2046 is 51 years
---
A(51) = 5.7*1.02^51 = 15.65 billion
------------------------
-In what year will the worlds population have tripled?
Ans:: When 1.02^x = 3
x = log(3)/log(1.02) = 55.48 years or the year 1995 + 55.48 = 2050
---------------------
-The predicted sustainable population of the earth is 20.1 billion people. When will Earth have too many people?
Solve:: 20.1 = 5.7(1.02)^x
3.52 = 1.02^x
x = log(3.52)/log(1.02) = 63.55 or the year 2058
-----------
Cheers,
Stan H.
---------
I got 15.6 billion for part a and I need help on parts b and c. Thank you!!